A 0.5 kg water balloon is dropped from a balcony 40 meters from the ground. What is the balloon's velocity just as it hits the ground? (Ignoring air resistance)
1
Expert's answer
2020-04-03T13:24:06-0400

Ignoring the air resistance means the case of free fall (gravity is the only force acting upon body).

"h=v_0 t+\frac{gt^2}{2}"

and

"v=v_0+gt",

where "h" - height, m; "v" - velocity, m/s; "v_0" - initial velocity, m/s; "t" - time, s; "g" - acceleration due to gravity, m/s2 ("g \approx 9.8 \; m/s^2"). In our case "v_0 =0", so balloon will hit the ground at

"t= \sqrt{\frac{2h}{g}}"

"v=g \cdot \sqrt{\frac{2h}{g}} = \sqrt{2gh}"

"v=\sqrt{2 \cdot 9.8 \cdot 40} = \sqrt{784} = 28 \; m/s"


Answer: 28 m/s.